OpenStudy (anonymous):

what is the radius of this equation: 3x squared plus 3y squared plus 12x minus 24y minus 48=0

OpenStudy (anonymous):

3x²+3y²+12x-24y-48=0 x²+y²+4x-8y-16=0 (x+2)²+(y-4)² =36 sqrt(36)=6 So the radius is 6.

OpenStudy (anonymous):

what is the radius of this equation: x squared plus y squared minus 14x plus 6y plus 42=0